Find the inverse Laplace transform of the following (i) 𝑠 (𝑠+π‘Ž) 2+𝑏2, (ii) 𝑠 2βˆ’5 𝑠 3+4𝑠 2+3𝑠 ’ (iii) 3𝑠 (𝑠+1) 4, (iv) 1 (𝑠 2+1) 2.Β 


1
Expert's answer
2021-09-29T08:59:44-0400

(i)


"f(t)=L^{-1}\{F(s)\}"

"=L^{-1}\{\dfrac{s}{(s+a)^2+b^2}\}=L^{-1}\{\dfrac{s+a-a}{(s+a)^2+b^2}\}"

"=L^{-1}\{\dfrac{s+a}{(s+a)^2+b^2}\}-L^{-1}\{\dfrac{a}{(s+a)^2+b^2}\}"

"=e^{-at}\cos(bt)-\dfrac{a}{b}e^{-at}\sin(bt)"

(ii)


"f(t)=L^{-1}\{F(s)\}=L^{-1}\{\dfrac{s^2-5}{s^3+4s^2+3s}\}"

"s^3+4s^2+3s=s(s^2+4s+3)=s(s+1)(s+3)"

"\dfrac{s^2-5}{s^3+4s^2+3s}=\dfrac{A}{s}+\dfrac{B}{s+1}+\dfrac{C}{s+3}"

"=\dfrac{A(s+1)(s+3)+Bs(s+3)+Cs(s+1)}{s^3+4s^2+3s}"

"s=0: -5=3A=>A=-\dfrac{5}{3}"

"s=-1: -4=-2B=>B=2"

"s=-3: 4=6C=>C=\dfrac{2}{3}"

"f(t)=L^{-1}\{\dfrac{s^2-5}{s^3+4s^2+3s}\}"

"=-\dfrac{5}{3}L^{-1}\{\dfrac{1}{s}\}+2L^{-1}\{\dfrac{1}{s+1}\}+\dfrac{2}{3}L^{-1}\{\dfrac{1}{s+3}\}"

"=-\dfrac{5}{3}+2e^{-t}+\dfrac{2}{3}e^{-3t}"

(iii)


"f(t)=L^{-1}\{F(s)\}"

"=L^{-1}\{\dfrac{3s}{(s+1)^4}\}=3L^{-1}\{\dfrac{s+1-1}{(s+1)^4}\}"

"=3L^{-1}\{\dfrac{1}{(s+1)^3}\}-3L^{-1}\{\dfrac{1}{(s+1)^4}\}"

"=\dfrac{3}{2}t^2e^{-t}-\dfrac{1}{2}t^3e^{-at}"

(iv)


"f(t)=L^{-1}\{F(s)\}=L^{-1}\{\dfrac{1}{(s^2+1)^2}\}"

"=\dfrac{1}{2(1)^3}(\sin(1\cdot t)-1\cdot t\cos(1\cdot t))"

"=\dfrac{1}{2}\sin(t)-\dfrac{1}{2}t\cos(t)"